OpenStudy (anonymous):

whats the volume of a square pyramid with a base of 16x16 and a slant height of 15

OpenStudy (anonymous):

o and my answer choices are 80, 240, 1082.8,and 1450.7

OpenStudy (anonymous):

hoblos - dont you use the perpenducular height in that formula? if my memory serves me right!

OpenStudy (hoblos):

15^2 = 8^2 +h^2 h=12.69 V=(1/3)(16^2)(12.69)=1082.8 @joewhit yeah you are right sorry for my first answer

OpenStudy (anonymous):

thanks!!!!

OpenStudy (hoblos):

yw :D